OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal+fan=you for BEST response. How much is a 15% tip for a meal costing $32? A. $4.80 B. $4.70 C. $2.13 D. $1.50

OpenStudy (rainbow_rocks03):

A?

OpenStudy (anonymous):

to say 15% of something, is same as saying \(\huge\frac{15}{100}\) of the same thing. So: \((32\times \large \frac{15}{100}) = \) ?

OpenStudy (rainbow_rocks03):

4.8

OpenStudy (anonymous):

right

OpenStudy (rainbow_rocks03):

So I was correct @myko ?

OpenStudy (anonymous):

yes

OpenStudy (rainbow_rocks03):

TY for the help but will you help me with 2 more questions plz?

OpenStudy (rainbow_rocks03):

@myko ?